OpenStudy (anonymous):

Ok, What am I doing wrong: 1/ (x^2 - 7x +10) = x/ (x-5) + 1/ (x-2) 1/ (x-5)(x-2) = x/ (x-5) + 1/ (x-2) LCM: (x-5)(x-2) so (x-5)(x-2)( 1/ (x-5)(x-2) = (x-5)(x-2)(x/ (x-5) + 1/ (x-2) 1= x(x-5) + 1(x-2) 1=x^2 - 5x +1x - 2 1= x^2 - 4x- 2 .... solve for x by factoring... ??? I know there's going to be two answers but I don't know what I'm doing wrong

jimthompson5910 (jim_thompson5910):

Your left side is correct

jimthompson5910 (jim_thompson5910):

But the right side is not correct

jimthompson5910 (jim_thompson5910):

what is x/ (x-5) multiplied by (x-5)(x-2) ???

OpenStudy (anonymous):

I have to multiply both sides by the LCM, simplify, then solve

jimthompson5910 (jim_thompson5910):

Yes and after you multiply the right side by (x-5)(x-2), you distribute that through

OpenStudy (anonymous):

so 1= x(x-2) + 1(x-5)? 1= x^2 -2x+1x-5 1=x^2-x-5?

jimthompson5910 (jim_thompson5910):

better

OpenStudy (anonymous):

then make is equal to zero so 0= x^2-x-6?

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

how would i factor that though?

jimthompson5910 (jim_thompson5910):

Find two numbers that multiply to -6 and add to -1

jimthompson5910 (jim_thompson5910):

Those two numbers are...???

OpenStudy (anonymous):

-3 and 2

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

So x^2-x-6 factors to (x-3)(x+2)

jimthompson5910 (jim_thompson5910):

which means x^2-x-6 = 0 is the same as (x-3)(x+2) = 0

OpenStudy (anonymous):

yay! ok.. thanks! x= -2 or 3

jimthompson5910 (jim_thompson5910):

perfect

OpenStudy (anonymous):

thank you!

jimthompson5910 (jim_thompson5910):

you're welcome